Paris Company purchased equipment on January 1, 2012, for $25,000. The estimated useful life of the equipment is five years, the salvage value is $5,000, and the company uses the double-declining-balance method to depreciate fixed assets. a. Provide the journal entry assuming the equipment is scrapped after three years. . Provide the journal entry assuming the equipment is scrapped after five years. Provide the journal entry assuming the equipment is sold for $8,000 after three years. . Provide the journal entry assuming, at the end of the fifth year, the equipment and $28,000 cash are traded in for a dissimilar asset with an objectively determined FMV of $30,000.
Answer #1
Depreciation by double declining method = 2 × Straight-line depreciation rate × Book value at the beginning of the year
Depreciation by staight line = 25000-5000(scrap value )/5 years= 20000/5= 4000
as %age this is 4000/20000= 20% there for %age for double declining will be 2 * 20 % = 40%
Depreciation table:
Depreciation Depreciation Accumulated Book
Date % Expense Cost Depreciation Value
01-01-12 $25,000 0 $25,000
12/31/12 40% $10,000 25,000 10,000 15,000
12/31/13 40% 6,000 25,000 16,000 9,000
12/31/14 40% 3,600 25,000 19,600 5,400
12/31/15 40% 400 25,000 20,000 5,000 equipment's book value cannot go below its estimated salvage value, depreciation expense for 2015 cannot exceed $400.
12/31/16 40% 0 25,000 20,000 5,000
Debit Credit
a. Accumulated Depreciation—Equipment 19,600
Loss on Disposal of Equipment 5,400
Equipment 25,000
(sale of equipment as scrap)
b. Accumulated Depreciation—Equipment 20,000
Loss on Disposal of Equipment 5,000
Equipment 25,000
(sale of equipment as scrap)
c. Cash 8,000
Accumulated Depreciation—Equipment 19,600
Equipment 25,000
Gain on Sale of Fixed Assets 2,600
Sold equipment for cash
d. Fixed Asset - new 30000
Accumulated Depreciation—Equipment 20,000
Loss on Disposal of Fixed Asset 3,000
Cash 28,000
Equipment 25000
Exchanged fixed asset for a dissimilar asset